An investment project provides cash inflows of $710 per year for 10 years.
Required:
(a) What is the project payback period if the initial cost is 2,130?
(Click to select)3.06 years2.91 years2.85 years3.03 years3 years

  

(b) What is the project payback period if the initial cost is 4,828?
(Click to select)6.8 years6.6 years6.87 years6.46 years6.94 years

  

(c) What is the project payback period if the initial cost is 7,810?
(Click to select)5.25 years3.01 yearsNever1.35 years4.95 years
Answer #1

payback period= initial cost/annual cash flows

1.payback period=2,130/710=3 years

2.payback period=4828/710=6.8 years

3.payback period=7810/710=Never.